Question 12 A bullet is fired horizontally from a handgun at a target 100.0 m away. If the initial speed of the bullet as it leaves the gun is 300.0 m/s, how far vertically will the bullet have dropped by the time it hits the target? Ignore the effects of air resistance NOTE TO PHYS 1A STUDENTS: This problem originally had an error, so it will not be counted in the HW score whether or not you already entered an answer, as some students had already used up their 2 attempts to enter answers. The error has now been corrected by the author, and the problem is still worth trying as a practice problem.] Select the correct answer CHECK ANSWER 0 of 1 attempts used LAST ATTEMPT O 3.1 m O 1.2 m O 1.1 m
Answer #1

Speed of bullet = 300 m/s

Target distance = 100 m

Total time taken = Distance/ speed

T = 100/300

T = 1/3 second

Now consider vertical motion of bullet.

Initial velocity of bullet in vertical direction = 0 m/s

Using equation

S0.5gt2

,S = 0 * t + 0.5 * q * t.

S = 0.54771

None of the option is correct.